OpenStudy (anonymous):

Solve 2x – 5 = 3(2x + 5).

OpenStudy (anonymous):

First expand teh brackets. Are you able to do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So: 2x-5 = (3 * 2x) + (3*5) 2x-5= 6x + 15

OpenStudy (anonymous):

Move the term with the lower x (2x is lower than 6x) to the other side by subtracting 2x on both sides.

OpenStudy (anonymous):

-5=6x - 2x +15 Simplify: -5=4x +15 -20=4x x=-5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you understand what I did?

OpenStudy (anonymous):

yes but is there two solutions?

OpenStudy (anonymous):

Hmm...I don't think so...why? Is there another solution?

OpenStudy (anonymous):

oh i am mistaken. sorry. thanks

OpenStudy (anonymous):

That's alright. Got another question you're stuck on?

OpenStudy (anonymous):

i will post it when i get stuck

OpenStudy (anonymous):

Ok :)